OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. 5 the square root of 2 + the square root of 2 10 6 the square root of 2 5 + the square root of 2 10 + the square root of 2

OpenStudy (cwrw238):

hint: 5x + x = 6x

OpenStudy (anonymous):

B?

OpenStudy (cwrw238):

treat sqrt 2 as x

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

thank you can you help me with more?

OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. the square root of 169 − the square root of 50 − the square root of 8 13 − 7 the square root of 2 13 − 3 the square root of 2 7 the square root of 2 6

OpenStudy (anonymous):

13 - 5 sqrt(2) - 2sqrt(2)

OpenStudy (cwrw238):

169 has an integer square root sqrt 50 = sqrt ( 25 * 2) = sqrt 25 * sqrt 2 sqrt 8 = sqrt 4 * sqrt 2

OpenStudy (anonymous):

thank you!

OpenStudy (cwrw238):

ok

OpenStudy (anonymous):

I think I understand it now